OpenStudy (anonymous):

Converting negative exponents to radicals: Will you see if I got this right?

OpenStudy (amistre64):

\[\left(\frac{a}{b}\right)^{-k}=\left(\frac{b}{a}\right)^{k}\]

OpenStudy (anonymous):

Is \[f'(x)=\frac{ 3 }{ 2 }x ^{\frac{ 1}{ 2 }}+\frac{ 1 }{ 2 }x ^{\frac{ -1 }{ 2 }}+2\] The same as: \[f'(x)=\frac{ 3\sqrt{x} }{ 2 }+\frac{ 1 }{ 2\sqrt{x} }+2\] ?

OpenStudy (anonymous):

yep. all good here

OpenStudy (amistre64):

1/2 exponent is sqrt, it looks fine to me